OpenStudy (anonymous):

Anyone know how to do these 3 problems ?? I got them wrong

OpenStudy (anonymous):

and this goes with it: Keng adds a 3-inch-wide frame around all sides of his canvas. Write a polynomial expression, in simplified form, that represents the total area of the canvas and the frame. Be sure to show your work. Keng is unhappy with his 3-inch-wide frame, so he decides to put a frame with a different width around his canvas. The total area of the canvas and the new frame is given by the polynomial h2 + 8h + 12, where h represents the height of the canvas. Determine the width of the new frame. Show all your work. Explain why you did each step.

OpenStudy (anonymous):

1. A= width*height= \[h^2 = (3)^2 = 9\] 2. A=width*height height is still going to be h Area is given in the problem (width)h = \[h^2 + 8h + 12\] After long division: \[h \sqrt{h^2}+8h+12\]= \[h+8+12/h\] -(h^2) 8h+12 -(8h) 12 3. A=width*height=h(h+4)=h^2+4h

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yep